OpenStudy (anonymous):

David has $2.25 consisting of 15 coins, all dimes and quarters, in his pocket. How many dimes and quarters does he have?

OpenStudy (anonymous):

10 dimes and 5 quarters I believe

OpenStudy (gw2011):

Let D=Dimes Let Q=Quarters D+Q=15 0.10D+0.25Q=2.25 Multiply the first equation by 0.25 to eliminate Q 0.25D+0.25Q=3.75 0.10D+0.25Q=2.25 Subtract this equation from the first one 0.15D=1.5 D=10 dimes D+Q=15 Q=15-10=5 Quarters Therefore, there are 10 dimes and 5 quarters